OpenStudy (briggles1999):

MEDAL AND FAN <3 Given the proportion a/b = 8/15, which ratio completes the equivalent proportion a/8 = ?

OpenStudy (briggles1999):

@AloneS

OpenStudy (briggles1999):

@math_man21 please help lol im stuck

OpenStudy (math_man21):

ok sure

OpenStudy (briggles1999):

Thank you <3

OpenStudy (math_man21):

what grade is this

OpenStudy (briggles1999):

10th

OpenStudy (math_man21):

sweet in 8th but do 10 work

OpenStudy (briggles1999):

awesome that must mean you are super smart lol

OpenStudy (math_man21):

yep

OpenStudy (briggles1999):

lol well this is an essay question and im not sure if im supposed to cross multiply or not

OpenStudy (math_man21):

id say so from looks of it

OpenStudy (briggles1999):

Can you help me explain this... i just dont grasp it

OpenStudy (math_man21):

a/b = 8/15, a/8 = ? so take that *(plus what u no)

OpenStudy (math_man21):

that works to alones

OpenStudy (math_man21):

*skakes hand and walkls away*

alones (alones):

No i'm askign if the problem looks like that

OpenStudy (briggles1999):

yes it looks like that and no i dont

alones (alones):

Oh okay..@math_man21 you can go on xP

OpenStudy (briggles1999):

im starting to wonder if its b/15... if a/8... than wouldnt it be b/15?